OpenStudy (anonymous):

To Applekiwi & anilorap. I tried to post an answer but it is hung up. The answer was: -3/4x - 17/2. So what did we do wrong?

OpenStudy (anonymous):

question?

OpenStudy (anonymous):

yup, the slope is -3/4, I've posted it.

OpenStudy (anonymous):

hang on, let me re-type the whole thing

OpenStudy (anonymous):

slope = {-4-(-7)} / {-6-(-2)} = 3/(-4) = -3/4 you have two method to find the y-int. I'll stick to the one posted on your other post: y-(-7) = -3/4(x- (-2)) y+7=-3/4(x+2) y = (-3/4)x - 3/2 - 7 y = (-3/4)x - 17/2